Given that x is a normal variable with mean μ = 51 and standard deviation σ = 6.1, find the following probabilities. (Round your answers to four decimal places.)

(a)  P(x ≤ 60)


(b)  P(x ≥ 50)


(c)  P(50 ≤ x ≤ 60)
Answer #1

Ans:

a)

z=(60-51)/6.1

z=1.475

P(z<=1.475)=0.9299

b)

z=(50-51)/6.1=-0.164

P(z>=-0.164)=0.5651

c)

P(-0.164<=z<1.475)=P(z<=1.475)-P(z<=-0.1640)

=0.9299-0.4349=0.4950